OpenStudy (anonymous):

Ok honestly the notes make no sense on this one... How do I do this? \[\huge \sum_{n=1}^{\infty} \frac{\sin(\frac{\pi n}{2})}{n!}\]

OpenStudy (anonymous):

I got to go for now but I'll be back tomorrow to see if anybody has figured this one out :-)

OpenStudy (anonymous):

Convergence/Divergence and if it converges to what are the typical questions for these...

OpenStudy (anonymous):

hi agentx5 1.Convergence/Divergence u can write \( \large 0 \le |\frac{\sin \frac{n \pi}{2}}{n!}| \le \frac{1}{n!} \) and use this \(\large \sum_{n=1}^{\infty} \frac{1}{n!}=e-1\) ------------------------------------------------------------------------- 2.Convergence to What see \(\large \sin \frac{n \pi}{2}=0 \ \ for \ \ n=2,4,6,... \) and \(\large \sin \frac{n \pi}{2}=1 \ or \ -1 \ \ for \ \ n=1,3,5,... \) so u can rearrange the sigma like this \[\large \sum_{n=0}^{\infty} \frac{(-1)^n}{(2n+1)!}\] now think of Maclaurin series for sin to Evaluate the sum

OpenStudy (anonymous):

So kind of like the Squeeze Theorem for the first one, gotcha. Does this test have a name though? Typically we have to say something like "convergent by the alternating series test" or something like that. I don't think we've gotten to MacLaurin series yet, but I'll try to understand :-) Would #46 from this be what I'm looking for? http://mathworld.wolfram.com/MaclaurinSeries.html

OpenStudy (anonymous):